Andy is driving at a speed of 95 MPH. How many miles has he traveled in 24 mins?
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(57387)

Andy is driving at a speed of 95 MPH. How many miles has he traveled in 24 mins?
-----
distance = rate*time
dist = 95(24/60 hr)
= 95(2/5)
= 19*2
= 38 miles
=================
cheers,
Stan H.

Answer by ewatrrr(10682)


Hi
D = r*t |Note: both rate and time must have same units
24min = 24/60 hr = 2/5 hr
D = 95mph*(2/5)hr
D = 38mi